# Unleashing Large Multimodal Model’s Ability for Painting 3D Objects with Realistic Materials

### v2

@@ Abstract
Physically realistic materials are pivotal in augmenting the realism of 3D assets across various applications and lighting conditions. However, existing 3D assets and generative models often lack authentic material properties. Manual assignment of materials using graphic software is a tedious and time-consuming task. In this paper, we exploit advancements in Multimodal Large Language Models (MLLMs), particularly GPT-4V, to present a novel approach, Make-it-Real: 1) We demonstrate that GPT-4V can effectively recognize and describe materials, allowing the construction of a detailed material library. 2) Utilizing a combination of visual cues and hierarchical text prompts, GPT-4V precisely identifies and aligns materials with the corresponding components of 3D objects. 3) The correctly matched materials are then meticulously applied as reference for the new SVBRDF material generation according to the original diffuse map, significantly enhancing their visual authenticity. Make-it-Real offers a streamlined integration into the 3D content creation workflow, showcasing its utility as an essential tool for developers of 3D assets. Our project website is at: https://SunzeY.github.io/Make-it-Real/.
Keywords:
Texture Synthesis, MLLMs, Material Refinement, 3D Content Creation
Figure 1: Make-it-Real can refine any diffuse-map-only 3D object from both CAD design and generative models.
@@ 1 Introduction
Traditional computer graphics methods have been either manually creating materials or reconstructing them from physical measurements. Emerging text-to-3D generative models [8, 30, 50, 6, 10, 18, 45] are successful in creating complex geometries and detailed appearances, but they struggle to generate physically realistic materials, hampering their practical applicability. Recent studies also consider advanced aspects of appearance generation [11, 60, 57, 9], although they use simplified material models, limiting the diversity and realism of the generated materials. Moreover, they often require prohibitive training time.
@@
The emergence of Multimodal Large Language Models (MLLMs) [5, 41, 12, 2, 21, 51] provides novel approaches to problem-solving. These models have powerful visual understanding and recognition capabilities, along with a vast repository of prior knowledge, which covers the task of material estimation, as shown in Fig. 2. Specifically, we are using GPT-4V(ision) for the identification and matching of materials. Specifically, we first create highly detailed descriptions for materials and build a comprehensive material library . Next, we use GPT-4V to identify and match materials for each part of the object, utilizing visual prompts[59] and hierarchical text prompts. Finally, we achieve precise transferring of materials from the material library to man-made meshes, text-to-3D models, and image-to-3D model generated meshes, endowing them with material properties.
Our work differs from the aforementioned studies by leveraging prior knowledge from foundation models like GPT-4V to help acquire materials, and leveraging existing libraries as references to create new materials. A direct application of our work is empowering mature 3D content generation models with enhanced realism, depth effects, and improved visual quality. Furthermore, this can greatly streamline the creation process for 3D asset developers, as they only need to paint the diffuse texture on the mesh, and then the material properties for each part can be automatically generated.
To achieve our objective of transforming rough, unrealistic 3D objects into ones with realistic materials, we address the problem in three steps: Firstly, we prepare a high-quality, selectable library of materials featuring detailed descriptions created by GPT-4V. This material library is constructed based on the MatSynth dataset, which comprises over 4,000 high-resolution, tileable materials available under the CC0 license. Secondly, to segment the 3D mesh into distinct components, we utilize segmentation and back-projection techniques on multi-view images to accurately segment texture maps according to different parts of the 3D object. Thirdly, suitable materials will be assigned to each part, which is the key part of this paper. It involves the accurate identification of the material for each component and matching it with the materials in the material library. Unlike previous material recognition methods[35, 4], the challenge lies in recognizing material from rendered images of 3D objects with only an diffuse map, which may be distorted, reflecting only the base color of the object (as shown in  Fig. 1). Also it may be affected by shading or coupled with the color of illumination, both can influence judgment. Therefore, the model requires a strong recognition ability for materials and prior knowledge of object types and material information. Finally, the selected materials are meticulously applied to the 3D assets, endowing them with a realistic visual quality.
@@
  • We present the first exploration of leveraging multimodal large language models (MLLMs), e.g., GPT-4V for material recognition and unleashing their potential in applying real-world materials to 3D objects.
@@
  • We develope an effective pipeline for texture segmentation, material identification and matching, enabling the high-quality application of materials to 3D assets.
@@ 2 Related Work
3D Object Generation. The generation of 3D models using deep learning methods has experienced rapid development in recent times. The mainstream research can be primarily divided into two categories. The first category relies on techniques that optimize a Neural Radiance Field (NeRF) [38] or 3D Gaussian [27] guided by 2D diffusion model through score-distillation-sampling(SDS) loss [43, 37, 34, 9, 54, 46, 49, 55] or Score Jacobian Chaining (SJC) [53]. The second category aims at obtaining 3D representation via direct inference model, e.g., Point-E [39], Shap-E [26], and LRM [23], proven fast with high quality through large-scale pretraining [22, 20, 48, 31, 58]. Although the capabilities of these methods are continuously improving, they still lack a high degree of realism in textures. More importantly, since the textures of 3D objects obtained by these methods are shaded, they cannot directly respond to lighting changes under different lighting conditions, leading to less realism. Although some works attempt to generate PBR textures, their results are considerably limited to generate physically realistic materials due to the sparse generative guidance of SDS [60, 9]. Our work is the first to introduce the prior knowledge of MLLMs into the texture synthesis process. Through extensive experiments, we verifies that our method can be seamlessly and effectively applied to generated 3D objects, facilitating their graphic downstream applications.
Material Capture and Generation. Recent studies such as MatPalette[35] and TwoshotBRDF[4] have made progress in the recognition and extraction of 3D object materials, allowing for the retrieval of SVBRDF information from images of real materials[35] and combining object shape and illumination [4]. On the other hand, works like Paint-it[60], Matlaber[57], Fantasia3D[9], and TANGO[11] focus on generating text-controllable 3D meshes with physically-based rendering material properties. However, these methods either require extensive training beforehand or suffer from long computational times for generating complete PBR maps due to the calculation of SDS Loss. Additionally, they are limited to synthesizing only a subset of PBR textures and cannot generate the full range of material properties, such as height and displacement.
@@
Figure 2: Motivative observation. This example illustrates that GPT-4V can identify and appropriately conjecture materials to object components.
@@ 3 Methodology
3.1 Problem Definition
The extraction and restoration of materials corresponding to 3D meshes represent a challenging problem. This complexity is particularly pronounced when identifying and segregating different material regions within 3D objects with texture map limited in diffuse RGB domain, lacking in realism and stereoscopic effects. The process necessitates accurately identifying the material type of each region on the 3D mesh based solely on its intrinsic color. Following this identification, an appropriate material from our library is selected to enhance the visual fidelity of the original mesh. Moreover, the process involves accurately mapping the generated 2D material maps back onto the 3D mesh, ensuring the preservation of the material’s natural appearance and physical properties.
@@ 3.1 Problem Definition
The specific problem definition is as follows: material maps primarily encompass seven attributes: Diffuse, Albedo, Normal, Roughness, Metalness, Height, and Specular, ensuring compatibility with both diffuse-based (e.g., [16]) and diffuse-based (e.g., [36]) workflows. We focus on the diffuse-based rendering for an easier application to the current 3D content generation model[60, 9]. Given the advancements in generating high-quality 3D shapes and diffuse maps, the restoration of realistic material properties remains a challenge. We highlight a novel problem: given a 3D mesh ($\tilde{O}$) and known diffuse ($\tilde{D}$) map, which reflect the object’s intrinsic appearance, the goal is to extract and restore all other SVBRDF attributes of the object.
$Make\text{-}it\text{-}Real(\tilde{O},\tilde{D})=\{N,R,M,H,S\}$ (1)
Our goal is to utilize the Make-it-Real pipeline for the identification and synthesis of Spatially Varing Bidirectional Reflectance Distribution (SVBRDF) material maps, leveraging the existing diffuse map and prior knowledge to query the most appropriate material attribute maps $SV=\{N,R,M,H,S\}$, to use as references before further refinement.
@@
Figure 3: Material Library with Fine-Grained Annotations. Utilizing GPT-4V model, we develop a material library, meticulously generating and cataloging comprehensive descriptions for each material. This structured repository facilitates hierarchical querying for material allocation in subsequent looking up processes.
3.2 Overview
Figure 4: Overall pipeline. This pipeline of Make-it-Real is composed of Multi-View Image Segmentation, MLLM-based Material Matching, and SVBRDF Maps Generation.
In this section, we introduce our material synthesis pipeline Make-it-Real. As depicted in  Fig. 4, the process begins with rendering images from three different perspectives of a 3D model. Subsequent steps involve using SemanticSAM [29] to semantically mask and annotate different parts of the model from varied viewpoints via Set-of-Mark (SoM) [59]. Following this, we employ the Multimodal Large Language Model (MLLM) to retrieve the best matched material to different parts. We partition the texture map with specific material type using back-projection. Then, we estimate the SVBRDF values by pixel, referenced by origianl diffuse and the queried material maps of different parts from our material library. ultimately yielding texture-aligned material maps of the object.
3.3 Material Library
We aim to leverage both the material priors of the mesh as predicted by GPT-4V and the characteristics of similar materials from existing materials used by artists as references. As a result, we construct a fine-grained, annotated material library for Make, illustrated in Fig. 3. The library comprises 1,394 unique, tileable materials, spanning 13 major types ranging from ceramics to wood. Each material is equipped with detailed textual annotations generated by GPT-4V, which precisely describe the material’s appearance characteristics, providing rich semantic information for subsequent material assignment processes.
Data Source. Our data primarily originates from the MatSynth dataset [52], which offers comprehensive material maps with 4K resolution under the CC0 license. Each material is represented by seven maps: diffuse, diffuse normal, height, roughness, metallic and specular. We select 13 major types of materials from this dataset and crawl the corresponding material sphere images. These images are subsequently utilized to query GPT-4V for generating detailed descriptions of each material.
Prompt Design. Our prompt design aims to enable GPT-4V to generate detailed descriptions of each material based on realistic material sphere images, focusing on distinguishing appearance features such as color, pattern, roughness, metalness, embossing pattern, and material condition, while avoiding irrelevant information like the overall shape of the object. Our goal is to generate highly content-relevant descriptions that not only focus on appearance and attributes but also include knowledge-related information to differentiate subtle differences between materials. This approach is intended to precisely capture and describe the characteristics of various materials, facilitating fine-grained retrieval by GPT-4V, rather than mere categorization.
By integrating a fine-grained, annotated PBR (Physically Based Rendering) material library with detailed textual annotations generated by GPT-4V, our approach provides rich and precise material descriptions for 3D meshes. This not only enhances the search efficiency and practicality of the material library but also offers crucial semantic information for subsequent material assignment processes, significantly improving the ability to extract and restore materials from rendered images of 3D meshes.
3.4 Multi-View Image Segmentation
To accurately identify and segment different material regions on 3D meshes with diffuse RGB, we adopt an innovative segmentation strategy based on multi-view 2D image rendering. First, we render a series of 2D images of the 3D mesh from different perspectives. The goal of this step is to capture a comprehensive view of the 3D model, ensuring that we can identify material regions from various angles. Subsequently, we employ the Semantic-SAM [29] model for precise segmentation of the 2D rendered images. To enhance the accuracy of the segmentation and reduce noise interference, we introduce a mask filtering processing module. This module effectively handles and reduces the overlap and conflict between different masks, ensuring the clarity and accuracy of the segmentation results.
3.5 MLLM-based Material Matching
Following the successful segmentation of regions for multi-views of 3D meshes, our subsequent objective focuses on the allocation of appropriate materials to these regions. To achieve this aim, we introduce a material retrieval strategy powered by Multimodal Large Language Models (MLLMs), specifically leveraging the visual cues proposed by the Set-of-Mark (SoM) [59] model in conjunction with the robust multimodal understanding capabilities of GPT-4V. To enhance the precision of recognition, we employ a meticulously designed combination of visual and hierarchical text prompts in tandem with GPT-4V to retrieve the most suitable materials, in Fig. 3. This approach signifies a novel integration of language model insights and visual perception for the advancement of material assignment in 3D modeling, underscoring the potential of MLLMs in facilitating complex, multimodal tasks in computer vision and graphics.
Visual Prompt Based on SoM. In our method, we draw inspiration from the Set-of-Mark (SoM) technique [59], which overlays an array of spatially coherent and semantically interpretable marks on images to stimulate the visual understanding capabilities of large multimodal models, such as GPT-4V. By incorporating SoM technology, we annotate each segmented region with a unique identifier as a visual cue, aiding the model in comprehending and analyzing each distinct material area. This approach enables a more nuanced interaction between the model and visual content, facilitating more effective and precise processing of material regions.
@@ 3.5 MLLM-based Material Matching
Hierarchical Text Prompt. To ensure efficiency and accuracy in assigning appropriate materials to segmented regions of 3D meshes, we adopt a hierarchical text prompting approach. This process initiates by identifying the primary material type corresponding to each region. Subsequently, hierarchical prompts guide GPT-4V in discerning specific subclasses within the primary material category. We retrieve all descriptions under this subclass material category to pinpoint the description that best matches the segmented block. Such hierarchical processing enables us to search our material library with greater granularity, identifying the most fitting description for each material segment. This approach not only facilitates the allocation of the most suitable materials to each area but also reduces memory and time expenditures.
3.6 SVBRDF Maps Generation
We conduct the generation of SVBRDF maps in two stages. Initially, preliminary texture map region segmentation is obtained through querying material regions. Subsequently, referenced by the original diffuse map of an object, an estimation of BRDF values in the pixel space is performed by querying the matched realistic material maps.
Region-level Texture Map Partitioning. After obtaining the partitioned masks for each view of the 3D object, combined with preliminary material assignments and the corresponding rendered images, we facilitate the extraction of local rendered images specific to masked regions. These localized images capture intricate details and textures associated with distinct material regions on the 3D model. Subsequently, by back-projecting these local rendered images onto the 3D mesh and employing UV mapping, we align masks on the 2D image with specific regions on the texture map, thereby segmenting the texture. This procedure delineates dedicated areas within the texture map for each material zone identified in the rendered images, meticulously preserving the uniqueness and characteristics of each material while ensuring precise correspondence and seamless integration across the 3D model. This culminates in an initial segmentation of the texture map, where each section correlates with the attributes of a precisely queried material sphere.
Pixel-level Diffuse-referenced Estimation. To precisely estimate the spatially varying BRDFs per pixel, we draw on the common practice among artists who typically use the diffuse map as a guide to create maps for other material properties when generating various texture maps. This process involves using the original object’s diffuse map as a reference, and the matched material sphere’s diffuse map as a key. We finds the nearest neighbor pixel index in the key diffuse for each pixel’s RGB value in the query diffuse pixel by pixel, which we employ the K-d tree algorithm to accelerate the querying process. We get SVBRDF values at the corresponding pixel location by querying the key material map. Notably, this procedure applies histogram equalization to both the query diffuse and key diffuse, normalizing them to a similar color space. This method enables the generation of a series of spatially variant BRDF maps, which maintain high consistency with the texture of the diffuse color.

### v3

# Make-it-Real: Unleashing Large Multimodal Model for Painting 3D Objects with Realistic Materials

@@ Abstract
Physically realistic materials are pivotal in augmenting the realism of 3D assets across various applications and lighting conditions. However, existing 3D assets and generative models often lack authentic material properties. Manual assignment of materials using graphic software is a tedious and time-consuming task. In this paper, we exploit advancements in Multimodal Large Language Models (MLLMs), particularly GPT-4V, to present a novel approach, Make-it-Real: 1) We demonstrate that GPT-4V can effectively recognize and describe materials. 2) Utilizing a combination of visual cues and hierarchical text prompts, GPT-4V precisely identifies and aligns materials with the corresponding components of 3D objects. 3) The correctly matched materials are then meticulously applied as reference for the new SVBRDF material generation according to the original albedo map, significantly enhancing their visual authenticity. Make-it-Real offers a streamlined integration into the 3D content creation workflow, showcasing its utility as an essential tool for developers of 3D assets. Our project website is at: https://SunzeY.github.io/Make-it-Real/.
Figure 1: Usage of Make-it-Real. Our method can refine a wide range of albedo-map-only 3D objects from both CAD design and generative models. Our method enhances the realism of objects, enables part-specific material assignment to objects and generate PBR maps that are compatible with downstream engines.
@@ 1 Introduction
Traditional computer graphics methods have been either manually creating materials or reconstructing them from physical measurements. Emerging text-to-3D generative models [8, 32, 6, 10, 18, 47] and image-to-3D generative models [63, 53, 24, 58, 21, 57] are successful in creating complex geometries and detailed appearances, but they struggle to generate physically realistic materials, hampering their practical applicability. Recent studies have also explored advanced aspects of appearance generation [11, 68, 64, 9, 38]. However, they often rely on simplified material models. For instance,  [38] lacks the ability to produce metallic maps. All these approaches do not generate corresponding displacement and height maps, which restricts the diversity and realism of the generated materials, especially regarding depth and tactile qualities. Furthermore, these methods typically require relatively long training and inference time. Considering the abundance of high-quality 3D assets online [15, 16] that lack material attributes, and the maturity of 3D generative models in geometry and albedo modeling, we aim to recover materials from high-quality geometry and base-colored 3D meshes.
However, extracting and recovering material representations for 3D meshes is challenging. Unlike previous material recognition methods [38, 4, 19], this difficulty is heightened when identifying and separating different material regions within 3D objects in constricted albedo textures. These maps only reflect the base color and can be distorted, as shown by the globe in the upper right corner of Figure 1. Additionally, shadows and lighting can affect judgment. Thus, the model must have strong material recognition capabilities and prior knowledge of object types and materials.
@@
The emergence of Multimodal Large Language Models (MLLMs) [5, 43, 12, 2, 22, 54] provides novel approaches to problem-solving. These models have powerful visual understanding and recognition capabilities, along with a vast repository of prior knowledge, which covers the task of material estimation. Specifically, we are using GPT-4V(ision) for matching of materials. We first create highly detailed descriptions for materials to build a comprehensive library. Next, we use GPT-4V to retrieve materials for each segmented part of the object, utilizing visual prompts[66] and hierarchical text prompts. Finally, we meticulously designed algorithms to generate SVBRDF maps with consistent albedo, achieving realistic visual quality.
Notably, our work differs from the aforementioned studies by leveraging prior knowledge from foundation models like GPT-4V to extract and infer materials in albedo-only constrained scenarios. Additionally, we utilize existing material libraries as references to generate corresponding SVBRDF maps on a designed region-to-pixel algorithm. As illustrated in Figure 1, our approach features: 1) Enhanced 3D mesh realism: Leveraging GPT-4V’s visual perception and external knowledge, our method improves the realism, depth, and visual quality of a wide range of mature 3D content generation models. 2) Part-specific material matching: Ensuring material consistency with a segmentation network and refinement process, enabling precise material property retrieval for each segment. 3) Rendering engine compatibility: Generating six comprehensive material maps(roughness, metallic, specular, normal, displacement, height), which are compatible with downstream rendering engines. Developers only need to paint albedo textures; material properties are then automatically generated, saving extensive time on detailed ambient occlusion masks and material map creation.
@@
  • We present the first exploration of leveraging multimodal large language models (MLLMs), e.g.,GPT-4V for material recognition and unleashing their potential in applying real-world materials to extensive 3D objects with albedo-only.
@@
  • We develope an effective pipeline for texture segmentation, material matching and SVBRDF maps generation, enabling the high-quality application of materials to 3D assets.
@@ 2 Related Work
3D Object Generation. The generation of 3D models using deep learning methods has experienced rapid development in recent times. The mainstream research can be primarily divided into two categories. The first category relies on techniques that optimize a Neural Radiance Field (NeRF) [40] or 3D Gaussian [29] guided by 2D diffusion model through score-distillation-sampling(SDS) loss [45, 39, 37, 9, 59, 49, 51, 61]. The second category aims at obtaining 3D representation via direct inference model, e.g.,, Point-E [41], Shap-E [27], and LRM [24], proven fast with high quality through large-scale pretraining [23, 21, 52, 33, 65, 53, 58, 60]. Although the capabilities of these methods are continuously improving, they still lack a high degree of realism in textures. More importantly, since the textures of 3D objects obtained by these methods are shaded, they cannot directly respond to lighting changes under different lighting conditions, leading to less realism. Although some works attempt to generate PBR textures, their results are considerably limited to generate physically realistic materials due to the low robustness of SDS [68, 9]. Our work is the first to introduce the prior knowledge of MLLMs into the texture synthesis process. We verifies that our method can be seamlessly and effectively applied to generated 3D objects, facilitating their downstream applications under different lighting condition.
Material Capture and Generation. Recent studies such as Material Palette[38], MatSim[19], TwoshotBRDF[4] have made progress in the recognition and extraction of 3D object materials, allowing for the retrieval of SVBRDF information from images of real materials[38] and combining object shape and illumination [4], but they fail to extract and infer the materials of 3D objects with only albedo. On the other hand, works like Paint-it[68], Matlaber[64], Collaborative[55], Fantasia3D[9], and TANGO[11] focus on generating text-controllable 3D meshes with physically-based rendering material properties. However, these methods require either extensive training time on BRDF datasets or long inference time for generating materials. Additionally, they are limited to synthesizing only a subset of PBR textures and cannot generate the full range of material properties, such as height and displacement, which are essential for the fine tactile perception of object surfaces.
@@
Figure 2: Overall pipeline. This pipeline of Make-it-Real is composed of image rendering and material segmentation, MLLM-based material retrieval, and SVBRDF Maps Generation. We finally use blender engine to conduct physically-based rendering.
@@ 3 Methodology
3.1 Preliminary
Physically Based Rendering (PBR) materials are a compact representation of the bi-directional reflectance distribution function (BRDF), which describes how light is reflected from a surface. PBR material maps primarily encompass seven attributes: Albedo, Roughness, Metallic, Normal, Specular, Height, and Displacement. Based on the rendering equation [28], given a location $x$ and the surface normal $n$, the incident light intensity at this point is denoted as $L_{i}(\omega_{i};x)$ along the direction $\omega_{i}$; BRDF $f_{r}(\omega_{o},\omega_{i};x)$ denotes the reflectance coefficient of the material viewing from direction $\omega_{o}$. The observed light intensity $L_{o}(\omega_{o};x)$ is calculated over the hemisphere $\Omega=\{\omega_{i}:\omega_{i}\cdot n>0\}$:
[…]
@@ 3.1 Preliminary
Given the advancements in generating high-quality 3D shapes with albedo maps, the restoration of realistic material properties remains a challenge. We highlight a novel problem: given a 3D mesh ($\tilde{O}$) and known Albedo ($\tilde{A}$) map, which reflect the object’s intrinsic appearance, the goal is to extract and restore all other SVBRDF attributes of the object, i.e. Make-it-Real($\tilde{O},\tilde{A}$) = $\{R,M,N,S,H,D\}$. Our setup supports the popular Cook-Torrance analytical BRDF model [13]. In this parameterization, the BRDF includes components for albedo $b_{a}\in\mathbb{R}^{3}$, metallic $b_{m}\in\mathbb{R}$, and roughness $b_{r}\in\mathbb{R}$. For more complex surface simulations, such as displacement and height modeling, we use the Blender rendering engine to simulate the BRDF function $f_{r}(\omega_{o},\omega_{i};x)$.
3.2 Make-it-Real: A Framework for Material Matching and Generataion
In this section, we outline our material matching and generation pipeline, illustrated in Figure 2, which encompasses three stages: rendering and segmenting 3D meshes, retrieving matching materials using MLLM, and generating spatially varying BRDF maps from coarse to fine.
3.2.1 Rendering and Material Segmentation
To accurately segment different material regions on 3D meshes with albedo maps, we propose an innovative segmentation strategy based on 2D image rendering in Figure 2 (a). Initially, we use rasterization to render the input albedo mesh from various viewpoints to obtain a series of images:
$\mathcal{I}(x,y)=\mathcal{R}\left(\text{UV}_{\text{map}}\left(\text{rasterize}(\tilde{O},v_{t},x,y)\right),\mathcal{T}\right)$ (2)
where $\mathcal{I}(x,y)$ is the pixel value at image coordinates $(x,y)$, $\text{rasterize}(\tilde{O},v_{t},\cdot)$ maps the 3D mesh $\tilde{O}$ from viewpoint $v_{t}$ to 2D screen coordinates, $\text{UV}_{\text{map}}(\cdot)$ converts rasterized coordinates to UV coordinates, $\mathcal{T}$ is the albedo map, and $\mathcal{R}(\cdot,\mathcal{T})$ samples color from the $\mathcal{T}$ using the UV coordinates.
For the rendered images, we employ the Semantic-SAM [31] to perform preliminary semantic segmentation. Empirically, we select the main viewpoint with the largest projected area of the mesh, as it is more likely to contain more details. To address potential over-segmentation, drawing inspiration from [67], we extract non-overlapping segments from the masks to form distinct patches, as detailed in our approach in Figure 4 (a). These patches are then merged based on similar colors to obtain the final material grouping. We incorporate Set-of-Mark [66] method to annotate each material segment with a unique identifier, sorted by area size from largest to smallest. This annotation acts as a visual cue, enhancing the visual comprehension capabilities of large multimodal language models.
@@ 3.2.1 Rendering and Material Segmentation
Figure 3: The process of MLLM retrieving materials from the Material Library. Utilizing GPT-4V model, we develop a material library, meticulously generating and cataloging comprehensive descriptions for each material. This structured repository facilitates hierarchical querying for material allocation in subsequent looking up processes.
3.2.2 MLLM-based Material Retrieval
Material library with fine-grained annotations.
To enable large multimodal language models to accurately retrieve and match materials, we construct a finely annotated material library, as shown in Figure 3. It is composed of three main components: comprehensive PBR texture maps, highly-detailed records, and a category tree. It comprises 1,400 unique, tileable materials spanning 13 primary categories and 80 subtypes. The data primarily derives from the [56], which offers comprehensive PBR material textures under a CC0 license with 4K resolution. Each material is represented by seven maps: albedo, roughness, metallic, specular, normal, displacement and height. Accompanying each material are highly-detailed annotations by GPT-4V, offering thorough descriptions of the material’s visual characteristics and rich semantic information for the subsequent retrieval process. Created by crawling material sphere images and constructing prompts, these annotations capture subtle differences between materials, facilitating precise retrieval by GPT-4V, as detailed in Section B.4.
Hierarchical prompting for material retrieval.
@@ 3.2.2 MLLM-based Material Retrieval
Due to the vast size of our material library, feeding all prompts to GPT-4V simultaneously proves inefficient and challenging for memory retention. To ensure efficient and accurate material allocation in segmented areas of 3D meshes, we adopt a hierarchical text prompting approach. The schematic of the designed prompt is shown on the left side of Figure 3, and a complete querying process unfolds in Section B.4. This method starts by identifying the primary material types corresponding to each labeled region. Subsequently, hierarchical prompts guide GPT-4V to distinguish specific subclasses within the main material categories. We retrieve all descriptions under these subclasses to ascertain the most fitting descriptions for the segmented blocks. This hierarchical processing enables a more granular search of our material library, identifying the optimal description for each material segment. This approach aids in assigning the most suitable materials to each region and reduces memory and time consumption.
3.2.3 SVBRDF Maps Generation
We propose a method to generate SVBRDF maps on a region-to-pixel scale in Figure 2 (c). Initially, we segment texture map in uv space based on queried material regions on 2D image space. We then estimate BRDF values in pixel space using the object’s original albedo map for reference, ensuring consistency with the albedo map. This approach effectively enhances the realism of rendered surfaces.
Region-level texture map partitioning.
Upon acquiring segmentation masks for material regions within the 2D rendered space of a 3D mesh, our objective transitions to transposing these segmentations from the 2D image space to the corresponding UV space. As described in Section 3.2.1,we extract 2D image features from 3D mesh points via rasterization, and then we apply the material masks to these features, facilitating the accurate transfer of segmentation to UV space. To project the image feature $I_{t}$ back to the texture atlas $\mathcal{T}_{t}$ with segmented image mask $m_{t}$ at the viewpoint $v_{t}$, we apply gradient-based optimization for $L_{t}$ over the values of $\mathcal{T}_{t}$ when rendered through the differential renderer $\mathcal{R}$, as presented in Equation 3. In Equation 4, we then compute the difference between $\mathcal{T}^{mask}_{t}$ and the initialized $\mathcal{T}_{t}$ to transfer the mask image feature $m_{t}$ into the texture space, represented by $m_{uv}$. The term $\sigma$ denotes the difference coefficient. Due to the limited perspectives available in rendering, we avoid using a naive median-filling approach to ensure that there are no missing areas on the texture map. Instead, we employ a block-centric clustering based on albedo, as illustrated in Figure 4 (b), to obtain cohesive and refined region masks. The process is shown in Section B.1
$\nabla_{\mathcal{T}_{t}}L_{t}=\left(\mathcal{R}(\text{mesh},\mathcal{T}_{t},v_{t})-I_{t}\right)\odot m_{t}\frac{\partial\mathcal{R}}{\partial\mathcal{T}_{t}}.$ (3)
$m_{uv}=\mathbf{1}\left(\sum_{i=1}^{3}|\mathcal{T}^{mask}_{t}-\mathcal{T}_{t}|>\sigma\right).$ (4)
Pixel-level albedo-referenced estimation.
To achieve precise estimation of spatially varying BRDF (SVBRDF) at the pixel level, we draw inspiration from techniques commonly utilized by artists in creating texture maps. Artists typically use albedo maps as a reference for constructing ambient occlusion masks and further generating SVBRDF maps for material properties, which occupies a significant time portion of appearance modeling. Our method involves using the albedo map of the original object as a reference to refine the retrieved materials. We enhance the querying process using a KD-Tree algorithm, which searches for the nearest neighbor pixel index in the key albedo(retrieved map) for each RGB value of the queried albedo(input map) pixel, detailed in Section B.2. This process ensures that areas with similar colors exhibit similar BRDF values, avoiding abrupt changes in material properties; For regions with greater color differences, the distribution of differences aligns consistently with the input albedo, to simulate variations such as embossments or scratches. We retrieves SVBRDF values at pixel level, maintaining texture consistency with the albedo and producing appropriate concave or smooth surface effects. We further analyze the effects in Section 4.2.
Figure 4: Illustrations of mask refinement in 2D image space and UV texture space. (a) We effectively cluster concise material-aware masks compared to original segmented parts from [31]. (b) We fix missing parts on the uv texture space to get a complete texture partition map.